# Source Code Summarization in the Era of Large Language Models

### v1

@@ Abstract
To support software developers in understanding and maintaining programs, various automatic (source) code summarization techniques have been proposed to generate a concise natural language summary (i.e., comment) for a given code snippet. Recently, the emergence of large language models (LLMs) has led to a great boost in the performance of code-related tasks. In this paper, we undertake a systematic and comprehensive study on code summarization in the era of LLMs, which covers multiple aspects involved in the workflow of LLM-based code summarization. Specifically, we begin by examining prevalent automated evaluation methods for assessing the quality of summaries generated by LLMs and find that the results of the GPT-4 evaluation method are most closely aligned with human evaluation. Then, we explore the effectiveness of five prompting techniques (zero-shot, few-shot, chain-of-thought, critique, and expert) in adapting LLMs to code summarization tasks. Contrary to expectations, advanced prompting techniques may not outperform simple zero-shot prompting. Next, we investigate the impact of LLMs’ model settings (including top_p and temperature parameters) on the quality of generated summaries. We find the impact of the two parameters on summary quality varies by the base LLM and programming language, but their impacts are similar. Moreover, we canvass LLMs’ abilities to summarize code snippets in distinct types of programming languages. The results reveal that LLMs perform suboptimally when summarizing code written in logic programming languages compared to other language types (e.g., procedural and object-oriented programming languages). Finally, we unexpectedly find that CodeLlama-Instruct with 7B parameters can outperform advanced GPT-4 in generating summaries describing code implementation details and asserting code properties. We hope that our findings can provide a comprehensive understanding of code summarization in the era of LLMs.
@@ I Introduction
Code comments are vital for enhancing program comprehension [1] and facilitating software maintenance [2]. While it is considered good programming practice to write high-quality comments, the process is often labor-intensive and time-consuming [2, 3, 4]. As a result, high-quality comments are frequently absent, mismatched, or outdated during software evolution, posing a common problem in the software industry [5, 6, 7, 8]. Automatic code summarization (or simply, code summarization), a hot research topic [9, 10, 11, 12], addresses this challenge by developing advanced techniques/models for automatically generating natural language summaries (i.e., comments) for code snippets, such as Java methods or Python functions, provided by developers.
Recently, with the success of large language models (LLMs) in natural language processing (NLP) [13, 14], an increasing number of software engineering (SE) researchers have started integrating them into the resolution process of various SE tasks [15, 16, 17, 18], such as code generation [19, 20], program repair [21, 22], and vulnerability detection/localization [23, 24]. In this study, we focus on the application of LLMs on the code summarization tasks. Figure 1 shows the general workflow of LLM-based code summarization and its effectiveness evaluation methods. In the summary generation process, the input consists of a code snippet and the expected summary category. The input is passed to a prompt generator equipped with various prompt engineering techniques (referred to as prompting technique), which constructs a prompt based on input. This prompt is then used to instruct LLMs to generate a summary of the expected type for the input code snippet. In the summary evaluation process, a common method used to automatically assess the quality of LLM-generated summaries is to compute the text or semantic similarity between the LLM-generated summaries and the reference summaries.
@@ II Background and Related Work
Code summarization is the task of automatically generating natural language summaries (also called comments) for code snippets. Such summaries serve various purposes, including but not limited to explaining the functionality of code snippets [41, 8, 28]. The research on code summarization can be traced back to as early as 2010 when Sonia Haiduc et al. [42] introduced automated text summarization technology to summarize source code. Later on, following the significant success of neural machine translation (NMT) research in the field of NLP [43, 44], a large number of researchers migrate its underlying encoder-decoder architecture to code summarization tasks [45, 46, 47, 9, 4, 12]. In the past two years, research on LLM-based code summarization has mushroomed. Fried et al. [48] introduce an LLM called InCoder, and try zero-shot training on the CodeXGLUE [49] Python dataset. InCoder achieves impressive results, but fine-tuned small PLMs like CodeT5 can still outperform the zero-shot setting. Ahmed et al. [25] investigate the effectiveness of few-shot prompting in adapting LLMs to code summarization and find that it can make Codex significantly outperform fine-tuned small PLMs (e.g., CodeT5). Given the concern of potential code asset leakage when using commercial LLMs (e.g., GPT-3.5), Su et al. [50] utilize knowledge distillation technology to distill small models from LLMs (e.g., GPT-3.5). Their experimental findings reveal that the distilled small models can achieve comparable code summarization performance to LLMs. Gao et al. [30] investigate the optimal settings for few-shot learning, including few-shot example selection methods, few-shot example order, and the number of few-shot examples. Geng et al. [28] investigate LLMs’ ability to address multi-intent comment generation. Ahmed et al. [51] propose to enhance few-shot samples with semantic facts automatically extracted from the source code. Sun et al. [27] design several heuristic questions to collect the feedback of ChatGPT, thereby finding an appropriate prompt to guide ChatGPT to generate in-distribution code summaries. Rukmono et al. [52] address the unreliability of LLMs in performing reasoning by applying a chain-of-thought prompting strategy. Recently, some studies [26, 53, 11] have also investigated the applicability of Parameter-Efficient Fine-Tuning (PEFT) techniques in code summarization tasks. In this paper, we focus on uncovering the effectiveness of various prompting techniques in adapting LLMs to code summarization without fine-tuning.
@@ III-B Experimental LLMs
CodeLlama-Instruct. Code Llama [56] is a family of LLMs for code based on Llama 2 [57]. It provides multiple flavors to cover a wide range of applications: foundation models, Python specializations (Code Llama-Python), and instruction-following models (Code Llama-Instruct) with 7B, 13B, and 34B parameters. Our study utilizes Code Llama-Instruct-7B.
StarChat-$\beta$. StarChat-$\beta$ [58] is an LLM with 16B parameters fine-tuned on StarCoderPlus [59]. Compared with StarCoderPlus, StarChat-$\beta$ excels in chat-based coding assistance.
@@
Model Settings. Apart from RQ3 where we investigate the impact of model settings, we uniformly set the temperature to 0.1 to minimize the randomness of LLM’s responses and highlight the impact of evaluation methods/prompting techniques/programming language types/summary categories.
@@ III-C Prompting Techniques
when asking Instruction 2, Instruction 1 and Response 1 are paired as history prompts and answers input into the LLM.
@@
Due to the page limit, we present examples of the aforementioned prompting techniques on our anonymous site [40].
@@ III-D Experimental Datasets
TABLE I: Datasets. PP: Procedural Programming Languages, OOP: Object-Oriented Programming Languages, SP: Scripting Programming Languages, FP: Functional Programming Languages, LP: Logic Programming Languages.
@@ IV-A RQ1: What evaluation methods are suitable for assessing the quality of summaries generated by LLMs?
iii. Methods based on summary-code semantic similarity assess the quality of the generated summary by computing the semantic similarity between the generated summary and the code snippet to be summarized. Unlike the first two methods, this type of evaluation method does not rely on reference summaries and can effectively avoid issues related to low-quality and outdated reference summaries. SIDE proposed by Mastropaolo et al. [10] is a representative of this type of method. It provides a continuous score ranging within [-1,1], where a higher value represents greater similarity. We present the scores reported by the above similarity-based evaluation methods in percentage.
@@
Human Evaluation Results. Table II shows the human evaluation scores for reference summaries and summaries generated by the four LLMs. Observe that the scores of reference summaries in the three datasets are between 3 and 3.5 points, suggesting that the quality of the reference summaries is not very high. Therefore, evaluation methods based on summary-summary/code similarity may not accurately assess the quality of LLM-generated summaries.
@@
Automated Evaluation Results. Table III displays the scores of the LLM-generated summaries reported by three categories of automated evaluation methods, and LLM-based evaluation methods. Observed that among the three methods based on summary-summary text similarity, 1) the BLEU-based and ROUGE-L-based methods give StarChat-$\beta$ the highest scores on all three datasets; 2) the METEOR-based method gives StarChat-$\beta$ the highest score (i.e., 18.19) on the Java dataset, while gives CodeLlama-Instruct the highest scores (i.e., 21.64 and 17.29) on the Python and C datasets. Among the four methods based on summary-summary semantic similarity, BERTScore, SBCS, and SBED give the best scores to StarChat-$\beta$, and USECS gives the best score of 50.69 to CodeLlama-Instruct on the Java dataset. On the Python and C datasets, the four methods consistently give the best scores to CodeLlama-Instruct and StarChat-$\beta$, respectively. The summary-code semantic similarity-based method SIDE gives the highest scores (i.e., 80.46 and 66.35) to StarChat-$\beta$ on the Java and C datasets, and the highest score (i.e., 41.71) to GPT-3.5 on the Python dataset. The four LLM-based methods consistently give the highest scores to GPT-4 on the Java and C datasets, while they consistently award the highest scores to GPT-3.5 on the Python dataset.
@@
TABLE III: Automated evaluation scores for reference and LLM-generated summaries. S-S Tex.Sim.: methods based on summary-summary text similarity; S-S Sem.Sim.: methods based on summary-summary semantic similarity; S-C Sem.Sim.: methods based on summary-code semantic similarity. We bold the best score in each column.
@@
BLEU METEOR ROUGE-L BERTScore SBCS SBED USECS SIDE CodeLlama-I StarChat-$\beta$ GPT-3.5 GPT-4
Java Reference / / / / / / / 86.15 1.42 2.58 3.08 2.8 3.19
CodeLlama-Instruct 13.00 17.90 32.21 87.94 59.61 86.88 50.69 46.62 2.32 2.80 3.28 3.64 3.93
StarChat-$\beta$ 18.95 18.19 38.43 88.69 61.97 83.45 50.57 80.46 2.24 1.94 2.42 2.50 3.18
GPT-3.5 12.49 16.74 31.87 87.73 59.47 88.11 48.87 62.04 2.40 2.40 3.72 3.82 4.00
GPT-4 9.46 17.02 28.36 86.72 58.83 89.27 46.50 36.12 2.44 2.60 4.10 4.50 4.17
Python Reference / / / / / / / 16.11 1.48 2.74 2.84 2.98 3.56
CodeLlama-Instruct 16.04 21.64 37.80 89.06 61.57 85.40 55.86 38.84 1.62 2.60 3.44 3.72 3.88
StarChat-$\beta$ 18.35 17.62 37.96 88.92 58.97 87.39 51.54 24.05 1.94 1.96 2.40 2.42 3.14
GPT-3.5 11.95 19.14 30.20 87.63 61.37 86.36 49.54 41.71 1.96 2.72 4.32 4.30 4.16
GPT-4 14.07 20.87 35.38 88.11 60.65 87.04 51.21 33.16 1.76 2.54 3.92 4.16 4.06
C Reference / / / / / / / 64.23 1.56 2.80 2.24 2.62 3.05
CodeLlama-Instruct 10.92 17.29 28.71 86.38 51.55 95.94 37.95 46.69 2.62 3.02 3.82 3.84 4.15
StarChat-$\beta$ 15.58 15.57 32.84 87.27 54.85 91.92 40.60 66.35 2.76 2.74 2.20 2.62 3.49
GPT-3.5 12.06 16.00 29.81 86.65 53.61 93.71 39.75 50.17 3.04 2.86 3.48 3.66 4.06
GPT-4 10.07 16.18 28.63 86.03 53.00 94.77 37.30 41.37 3.18 2.86 4.00 4.36 4.25
@@
Correlation between Automated Evaluation and Human Evaluation. From Table III, it can also be observed that the average scores of reference summaries evaluated by the four LLM-based methods are mostly below 3 points. It means that similar to human evaluation, LLM-based evaluation methods also believe that the quality of the reference summaries is not very high. Besides, LLM-based evaluation methods are inclined to give higher scores to general-purpose LLMs GPT-3.5 and GPT-4, which is the same as human evaluation.
@@
Based on the above observations, we can reasonably speculate that compared to methods based on summary-summary text/semantic similarity and summary-code semantic similarity, LLM-based evaluation methods may be more suitable for evaluating the quality of summaries generated by LLMs. Therefore, we follow [9, 73] and calculate Spearman’s correlation coefficient $\rho$ with the $p$-value between the results of each automated evaluation method and human evaluation, providing more convincing evidence for this speculation. The Spearman’s correlation coefficient $\rho\in[-1,1]$ is suitable for judging the correlation between two sequences of discrete ordinal/continuous data, with a higher value representing a stronger correlation [77]. $-1\leq\rho<0$, $\rho=0$, and $0<\rho\leq 1$ respectively indicate the presence of negative correlation, no correlation, and positive correlation [78]. The $p$-value helps determine whether the observed correlation is statistically significant or simply due to random chance. By comparing the $p$-value to a predefined significance level (typically 0.05), we can decide whether to reject the null hypothesis and conclude that the correlation is statistically significant. Table IV shows the statistical results of $\rho$ and $p$-value. It can be clearly observed that among all automated evaluation methods, there is a significant positive correlation ($0.28\leq\rho\leq 0.65$) between the GPT-4-based evaluation method and human evaluation in scoring the quality of summaries generated by most LLMs, followed by the GPT-3.5-based evaluation method. For other automated evaluation methods, in most cases, their correlation with human evaluation is negative or weakly positive. Based on the above observations, we draw the conclusion that compared with other automated evaluation methods, the GPT-4-based method is more suitable for evaluating the quality of summaries generated by LLMs. In the subsequent RQs, we uniformly employ the GPT-4-based method to assess the quality of LLM-generated summaries. To make the output scores of GPT-4 more deterministic, we set the temperature to 0 when using GPT-4 as the evaluator.
TABLE IV: Spearman’s correlation coefficient $\rho$ with the $p$-value (values in parentheses) between the results of each automated evaluation method and human evaluation. CodeLlama-I: CodeLlama-Instruct. We bold the best score in each row.
Language Summary from S-S Tex.Sim. S-S Sem.Sim. S-C Sem.Sim. LLM-based Evaluation Method
BLEU METEOR ROUGE-L BERTScore SBCS SBED USECS SIDE CodeLlama-I StarChat-$\beta$ GPT-3.5 GPT-4
Java Reference / / / / / / / 0.31 (.03) 0.06 (.70) 0.26 (.07) 0.53 (.00) 0.60 (.00)
CodeLlama-I -0.25 (.08) 0.07 (.65) -0.14 (.35) -0.16 (.27) 0.01 (.94) -0.01 (.94) 0.01 (.95) -0.30 (.03) -0.00 (.98) -0.02 (.91) 0.35 (.01) 0.49 (.00)
StarChat-$\beta$ 0.22 (.13) 0.32 (.02) 0.15 (.31) 0.22 (.12) 0.30 (.04) -0.30 (.04) 0.32 (.02) -0.12 (.42) 0.11 (.44) 0.16 (.27) 0.52 (.00) 0.41 (.00)
GPT-3.5 -0.10 (.50) 0.12 (.39) 0.22 (.13) 0.0 (.99) -0.01 (.97) 0.01 (.97) 0.10 (.50) -0.20 (.16) 0.28 (.05) -0.08 (.57) 0.60 (.00) 0.56 (.00)
GPT-4 0.05 (.74) -0.02 (.87) 0.05 (.75) 0.02 (.87) 0.14 (.34) -0.14 (.34) 0.11 (.46) -0.19 (.18) 0.00 (.98) 0.04 (.76) 0.38 (.01) 0.40 (.00)
Python Reference / / / / / / / -0.04 (.79) -0.15 (.30) 0.19 (.19) 0.35 (.01) 0.37 (.01)
CodeLlama-I -0.17 (.24) -0.05 (.76) -0.11 (.46) -0.16 (.28) -0.08 (.56) 0.08 (.56) -0.08 (.60) 0.10 (.48) 0.15 (.29) -0.01 (.97) 0.52 (.00) 0.45 (.00)
StarChat-$\beta$ 0.20 (.17) 0.52 (.00) 0.19 (.20) 0.24 (.1) 0.18 (.22) -0.18 (.22) 0.50 (.00) 0.02 (.89) -0.19 (.18) 0.24 (.09) 0.48 (.00) 0.48 (.00)
GPT-3.5 -0.20 (.17) 0.01 (.96) -0.18 (.21) -0.07 (.61) -0.02 (.89) 0.02 (.89) -0.12 (.42) 0.17 (.24) 0.19 (.19) 0.35 (.01) 0.32 (.02) 0.42 (.00)
GPT-4 -0.24 (.09) 0.12 (.41) -0.16 (.26) -0.05 (.73) 0.02 (.91) -0.02 (.91) -0.19 (.19) 0.15 (.31) 0.18 (.21) 0.17 (.25) 0.16 (.28) 0.17 (.25)
C Reference / / / / / / / 0.23 (.11) 0.12 (.39) 0.27 (.06) 0.60 (.00) 0.62 (.00)
CodeLlama-I -0.28 (.05) -0.11 (.46) -0.09 (.55) -0.28 (.05) -0.05 (.76) 0.04 (.76) -0.07 (.61) -0.32 (.03) -0.00 (.98) -0.01 (.48) 0.33 (.02) 0.42 (.00)
StarChat-$\beta$0.00 (.98) 0.35 (.01) 0.17 (.24) 0.0 (.98) 0.17 (.22) -0.17 (.22) 0.26 (.07) -0.01 (.94) 0.05 (.72) 0.17 (.23) 0.16 (.27) 0.62 (.00)
GPT-3.5 -0.35 (.01) 0.14 (.33) -0.12 (.41) -0.06 (.67) 0.05 (.71) -0.05 (.71) 0.07 (.64) -0.47 (.00) 0.21 (.15) 0.21 (.14) 0.38 (.01) 0.65 (.00)
GPT-4 0.11 (.44) 0.38 (.01) 0.37 (.01) 0.15 (.29) 0.21 (.14) -0.21 (.15) 0.21 (.14) -0.26 (.07) -0.09 (.55) 0.01 (.92) 0.38 (.01) 0.28 (.05)
@@
TABLE V: Effectiveness of different prompting techniques

### v2

@@ Abstract
To support software developers in understanding and maintaining programs, various automatic (source) code summarization techniques have been proposed to generate a concise natural language summary (i.e., comment) for a given code snippet. Recently, the emergence of large language models (LLMs) has led to a great boost in the performance of code-related tasks. In this paper, we undertake a systematic and comprehensive study on code summarization in the era of LLMs, which covers multiple aspects involved in the workflow of LLM-based code summarization. Specifically, we begin by examining prevalent automated evaluation methods for assessing the quality of summaries generated by LLMs and find that the results of the GPT-4 evaluation method are most closely aligned with human evaluation. Then, we explore the effectiveness of five prompting techniques (zero-shot, few-shot, chain-of-thought, critique, and expert) in adapting LLMs to code summarization tasks. Contrary to expectations, advanced prompting techniques may not outperform simple zero-shot prompting. Next, we investigate the impact of LLMs’ model settings (including top_p and temperature parameters) on the quality of generated summaries. We find the impact of the two parameters on summary quality varies by the base LLM and programming language, but their impacts are similar. Moreover, we canvass LLMs’ abilities to summarize code snippets in distinct types of programming languages. The results reveal that LLMs perform suboptimally when summarizing code written in logic programming languages compared to other language types (e.g., procedural and object-oriented programming languages). Finally, we unexpectedly find that CodeLlama-Instruct with 7B parameters can outperform advanced GPT-4 in generating summaries describing code design rationale and asserting code properties. We hope that our findings can provide a comprehensive understanding of code summarization in the era of LLMs.
@@ I Introduction
Code comments are vital for enhancing program comprehension [1] and facilitating software maintenance [2]. While it is considered good programming practice to write high-quality comments, the process is often labor-intensive and time-consuming [2, 3, 4]. As a result, high-quality comments are frequently absent, mismatched, or outdated during software evolution, posing a common problem in the software industry [5, 6, 7]. Automatic code summarization (or simply, code summarization), a hot research topic [8, 9, 10], addresses this challenge by developing advanced techniques/models for automatically generating natural language summaries (i.e., comments) for code snippets, such as Java methods or Python functions, provided by developers.
Recently, with the success of large language models (LLMs) in natural language processing (NLP) [11], an increasing number of software engineering (SE) researchers have started integrating them into the resolution process of various SE tasks [12, 13]. In this study, we focus on the application of LLMs on the code summarization tasks. Figure 1 shows the general workflow of LLM-based code summarization and its effectiveness evaluation methods. In the summary generation process, the input consists of a code snippet and the expected summary category. The input is passed to a prompt generator equipped with various prompt engineering techniques (referred to as prompting technique), which constructs a prompt based on input. This prompt is then used to instruct LLMs to generate a summary of the expected type for the input code snippet. In the summary evaluation process, a common method used to automatically assess the quality of LLM-generated summaries is to compute the text or semantic similarity between the LLM-generated summaries and the reference summaries.
@@ II Background and Related Work
Code summarization is the task of automatically generating natural language summaries (also called comments) for code snippets. Such summaries serve various purposes, including but not limited to explaining the functionality of code snippets [30, 7, 17]. The research on code summarization can be traced back to as early as 2010 when Sonia Haiduc et al. [31] introduced automated text summarization technology to summarize source code. Later on, following the significant success of neural machine translation (NMT) research in the field of NLP [32, 33], a large number of researchers migrate its underlying encoder-decoder architecture to code summarization tasks [34, 35, 36, 8, 4]. In the past two years, research on LLM-based code summarization has mushroomed. Fried et al. [37] introduce an LLM called InCoder, and try zero-shot training on the CodeXGLUE [38] Python dataset. InCoder achieves impressive results, but fine-tuned small PLMs like CodeT5 can still outperform the zero-shot setting. Ahmed et al. [14] investigate the effectiveness of few-shot prompting in adapting LLMs to code summarization and find that it can make Codex significantly outperform fine-tuned small PLMs (e.g., CodeT5). Given the concern of potential code asset leakage when using commercial LLMs (e.g., GPT-3.5), Su et al. [39] utilize knowledge distillation technology to distill small models from LLMs (e.g., GPT-3.5). Their experimental findings reveal that the distilled small models can achieve comparable code summarization performance to LLMs. Gao et al. [19] investigate the optimal settings for few-shot learning, including few-shot example selection methods, few-shot example order, and the number of few-shot examples. Geng et al. [17] investigate LLMs’ ability to address multi-intent comment generation. Ahmed et al. [40] propose to enhance few-shot samples with semantic facts automatically extracted from the source code. Sun et al. [16] design several heuristic questions to collect the feedback of ChatGPT, thereby finding an appropriate prompt to guide ChatGPT to generate in-distribution code summaries. Rukmono et al. [41] address the unreliability of LLMs in performing reasoning by applying a chain-of-thought prompting strategy. Recently, some studies [15, 42, 10] have also investigated the applicability of Parameter-Efficient Fine-Tuning (PEFT) techniques in code summarization tasks. In this paper, we focus on uncovering the effectiveness of various prompting techniques in adapting LLMs to code summarization without fine-tuning.
@@ III-B Experimental LLMs
CodeLlama-Instruct. Code Llama [45] is a family of LLMs for code based on Llama 2 [46]. It provides multiple flavors to cover a wide range of applications: foundation models, Python specializations (Code Llama-Python), and instruction-following models (Code Llama-Instruct) with 7B, 13B, 34B, and 70B parameters. In this study, we evaluate the summaries generated by CodeLlama-Instruct 7B. We also verify the ability of CodeLlama-Instruct 70B to act as an evaluator in RQ1.
$\beta. StarChat$-$\beta$ [47] is an LLM with 16B parameters fine-tuned on StarCoderPlus [48]. Compared with StarCoderPlus, StarChat-$\beta$ excels in chat-based coding assistance.
@@
Model Settings. Apart from RQ3 where we investigate the impact of model settings, we uniformly set the temperature to 0.1 to minimize the randomness of LLM’s responses and highlight the impact of evaluation methods/prompting techniques/programming language types/summary categories. Note that the LLM-generated summaries are relatively long, and usually consist of several sentences. By observing some examples, we find that the first sentences are suitable as the final summaries, and the following sentences elaborate on some details and supplementary explanations. Therefore, we extract the first sentences as the final summaries.
@@ III-C Prompting Techniques
when asking Instruction 2, Instruction 1 and Response 1 are paired as history prompts and answers and input into the LLM.
@@ III-D Experimental Datasets
TABLE I: Datasets. $P$: Procedural Programming Languages, OOP: Object-Oriented Programming Languages, SP: Scripting Programming Languages, FP: Functional Programming Languages, LP: Logic Programming Languages.
@@ IV-A RQ1: What evaluation methods are suitable for assessing the quality of summaries generated by LLMs?
iii. Methods based on summary-code semantic similarity assess the quality of the generated summary by computing the semantic similarity between the generated summary and the code snippet to be summarized. Unlike the first two methods, this type of evaluation method does not rely on reference summaries and can effectively avoid issues related to low-quality and outdated reference summaries. SIDE proposed by Mastropaolo et al. [9] is a representative of this type of method. It is based on contrastive learning and has been trained to assess the relevance of a given textual summary for a Java method. Note that it has not been trained on other language datasets. Hence, in our experiment, SIDE is only used to evaluate the Java dataset. SIDE provides a continuous score ranging within [-1,1], where a higher value represents greater similarity. We present the scores reported by the above similarity-based evaluation methods in percentage.
@@
Human Evaluation Results. Table II shows the human evaluation scores for reference summaries and summaries generated by the four LLMs. Observe that the scores of reference summaries in the three datasets are between 3 and 3.5 points, suggesting that the quality of the reference summaries is not very high. Therefore, evaluation methods based on summary-summary similarity may not accurately assess the quality of LLM-generated summaries.
@@
Automated Evaluation Results. Table III displays the scores of the LLM-generated summaries reported by three categories of automated evaluation methods, and LLM-based evaluation methods. Observed that among the three methods based on summary-summary text similarity, 1) the BLEU-based and ROUGE-L-based methods give StarChat-$\beta$ the highest scores on all three datasets; 2) the METEOR-based method gives StarChat-$\beta$ the highest score (i.e., 18.19) on the Java dataset, while gives CodeLlama-Instruct the highest scores (i.e., 21.64 and 17.29) on the Python and C datasets. Among the four methods based on summary-summary semantic similarity, BERTScore, SBCS, and SBED give the best scores to StarChat-$\beta$, and USECS gives the best score of 50.69 to CodeLlama-Instruct on the Java dataset. On the Python and C datasets, the four methods consistently give the best scores to CodeLlama-Instruct and StarChat-$\beta$, respectively. The summary-code semantic similarity-based method SIDE gives the highest score (i.e., 80.46) to StarChat-$\beta$ on the Java dataset. On the Java dataset, the five LLM-based methods consistently give the highest scores to GPT-4. While on the Python dataset, they consistently award the highest scores to GPT-3.5. On the C dataset, StarChat-$\beta$ gives the highest score to CodeLlama-Instruct while other LLMs give the highest score to GPT-4.
@@
TABLE III: Automated evaluation scores for reference and LLM-generated summaries. S-S Tex.Sim.: methods based on summary-summary text similarity; S-S Sem.Sim.: methods based on summary-summary semantic similarity; S-C Sem.Sim.: methods based on summary-code semantic similarity. CodeLlama-I: CodeLlama-Instruct. We bold the best score in each column.
@@
BLEU METEOR ROUGE-L BERTScore SBCS SBED USECS SIDE CodeLlama-I (7B) CodeLlama-I (70B) StarChat-$\beta$ GPT-3.5 GPT-4
Java Reference / / / / / / / 86.15 1.42 2.62 2.58 3.08 2.8 3.19
CodeLlama-I 13.00 17.90 32.21 87.94 59.61 86.88 50.69 46.62 2.32 3.72 2.80 3.28 3.64 3.93
StarChat-$\beta$ 18.95 18.19 38.43 88.69 61.97 83.45 50.57 80.46 2.24 2.52 1.94 2.42 2.50 3.18
GPT-3.5 12.49 16.74 31.87 87.73 59.47 88.11 48.87 62.04 2.40 3.38 2.40 3.72 3.82 4.00
GPT-4 9.46 17.02 28.36 86.72 58.83 89.27 46.50 36.12 2.44 3.88 2.60 4.10 4.50 4.17
Python Reference / / / / / / / / 1.48 3.13 2.74 2.84 2.98 3.56
CodeLlama-I 16.04 21.64 37.80 89.06 61.57 85.40 55.86 / 1.62 3.50 2.60 3.44 3.72 3.88
StarChat-$\beta$ 18.35 17.62 37.96 88.92 58.97 87.39 51.54 / 1.94 2.36 1.96 2.40 2.42 3.14
GPT-3.5 11.95 19.14 30.20 87.63 61.37 86.36 49.54 / 1.96 4.10 2.72 4.32 4.30 4.16
GPT-4 14.07 20.87 35.38 88.11 60.65 87.04 51.21 / 1.76 3.42 2.54 3.92 4.16 4.06
C Reference / / / / / / / / 1.56 2.25 2.80 2.24 2.62 3.05
CodeLlama-I 10.92 17.29 28.71 86.38 51.55 95.94 37.95 / 2.62 3.61 3.02 3.82 3.84 4.15
StarChat-$\beta$ 15.58 15.57 32.84 87.27 54.85 91.92 40.60 / 2.76 2.64 2.74 2.20 2.62 3.49
GPT-3.5 12.06 16.00 29.81 86.65 53.61 93.71 39.75 / 3.04 3.38 2.86 3.48 3.66 4.06
GPT-4 10.07 16.18 28.63 86.03 53.00 94.77 37.30 / 3.18 3.72 2.86 4.00 4.36 4.25
@@
Correlation between Automated Evaluation and Human Evaluation. From Table III, it can be observed that the average scores of reference summaries evaluated by the four LLM-based methods are mostly below 3 points. It means that similar to human evaluation, LLM-based evaluation methods also believe that the quality of the reference summaries is not very high. Besides, LLM-based evaluation methods are inclined to give higher scores to general-purpose LLMs GPT-3.5 and GPT-4, which is the same as human evaluation.
@@
Based on the above observations, we can reasonably speculate that compared to methods based on summary-summary text/semantic similarity and summary-code semantic similarity, LLM-based evaluation methods may be more suitable for evaluating the quality of summaries generated by LLMs. Therefore, we follow [8, 62] and calculate Spearman’s correlation coefficient $\rho$ with the $p$-value between the results of each automated evaluation method and human evaluation, providing more convincing evidence for this speculation. The Spearman’s correlation coefficient $\rho\in[-1,1]$ is suitable for judging the correlation between two sequences of discrete ordinal/continuous data, with a higher value representing a stronger correlation [66]. $1\leq\rho<0$, $\rho=0$, and $0<\rho\leq 1$ respectively indicate the presence of negative correlation, no correlation, and positive correlation [67]. The $p$-value helps determine whether the observed correlation is statistically significant or simply due to random chance. By comparing the $p$-value to a predefined significance level (typically 0.05), we can decide whether to reject the null hypothesis and conclude that the correlation is statistically significant. Due to the page limit, we present the statistical results of $\rho$ and $p$-value in [29]. The results demonstrate that among all automated evaluation methods, there is a significant positive correlation between the GPT-4-based evaluation method and human evaluation in scoring the quality of summaries generated by most LLMs, followed by the GPT-3.5-based evaluation method. For other automated evaluation methods, in most cases, their correlation with human evaluation is negative or weakly positive. In a shot, existing metrics are inadequate for LLM-generated summaries. The main reason for this phenomenon lies in the low quality of reference summaries. The reference summaries are short and detail-lacking, while LLMs can generate detailed useful summaries not fully covered by the reference summaries. Therefore, the similarity with the reference summary cannot accurately reflect the quality of the LLM-generated summaries. In addition, BLEU penalizes long summaries, which also affects the correlation between BLEU and human evaluation. Although SIDE does not rely on reference summaries, it is trained on reference summaries and is therefore also affected. Re-training SIDE on LLM-generated summaries requires high-quality and extensive training data, while collecting large-scale LLM-generated summaries, especially from LLMs like GPT-4, is costly. Numerous factors influence LLM-generated summary quality, complicating the collection of high-quality data. Based on the above observations, we draw the conclusion that compared with other automated evaluation methods, the GPT-4-based method is more suitable for evaluating the quality of summaries generated by LLMs. In the subsequent RQs, we uniformly employ the GPT-4-based method to assess the quality of LLM-generated summaries. We set the temperature to default value 1 when using GPT-4 as the evaluator.
[…]
Why do LLMs give such a score?
To investigate the scoring criteria of LLMs, we ask them to give explanations for their evaluation scores. The prompt we use is “Here is a piece of code with corresponding comments and your previous rate for each comment on a scale from 1 to 5, where a higher score indicates better quality. Please explain why you scored this way.
(a) A Java code snippet $c_{1}$
(b) Reference summary and LLM-generated summaries for $c_{1}$
Figure 3: An example of code, the reference summary, and LLM-generated summaries
(a) CodeLlama-Instruct (70B)’s explanations
(b) StarChat-$\beta$’s explanations
(c) GPT-3.5’s explanations
(d) GPT-4’s explanations
Figure 4: LLMs’ explanations of evaluation scores for the reference summary and LLM-generated summaries of $c_{1}$
Figure 4 shows LLMs’ explanations of evaluation scores for the reference summary and generated summaries of the example in Figure 3. Due to the page limit, we only show the results of CodeLlama-Instruct (70B), which performs better, instead of CodeLlama-Instruct (7B). Observe that LLM’s scoring criteria include accuracy, clarity, informativeness, and conciseness. Accuracy focuses on whether there is misleading information. Clarity focuses on whether the expression of a sentence is easy to understand. For example, the reference summary is evaluated by GPT-4 as not clear because “‘FROM REPLY - TO’ is not a standard way to refer to email fields”. Informativeness focuses on whether the summary contains enough details. As for conciseness, since the summaries generated by GPT-3.5 and GPT-4 are also evaluated as concise, we can infer that LLMs’ definition of concise is not equal to being as brief as the reference summary and the LLM-generated summaries are also considered concise. Therefore, the main factors affecting LLMs’ scoring are accuracy, clarity, and informativeness. That is, whether there is misleading information, whether the expression is easy to understand, and whether the summary contains enough details. This also explains why LLMs do not give high scores to reference summaries. Reference summaries are generally short, do not cover every detail in the code snippet, and may contain terms in the code snippet which increases the difficulty of understanding. However, although the scoring criteria of LLMs are similar, the scoring results are still different. This is because LLMs’ capabilities to understand code and summaries are different. As the SOTA LLM, GPT-4 achieves the best performance. It can accurately identify whether each summary contains all the details, while other LLMs may miss some of the details according to their explanation. In addition, GPT-4 does a better job in judging whether the summary is clear, as it points out that the phrase “To, Cc, and Bcc” is not easy to understand while GPT-3.5 and CodeLlama-Instruct (70B) does not point out this problem.
@@
TABLE IV: Effectiveness of different prompting techniques